Why Marginal Utility MUy of XY + 2X equals X


1
Expert's answer
2021-07-14T11:58:25-0400

Marginal utility =change in number of units consumed/change in total utility


Let q=Xy +2x

Taking MUy= dq/Dy of Xy+2x

Then MUy=x+0

=X